Any marks not recognised by the Optical Mark Recognition (OMR)

machine will get ZERO marks

Correct mistakes with an eraser, correction pen or white tape; it is not enough to put a cross through the

circle

Errors in student number get ZERO marks

Operating Systems and Systems Integration Quick Quiz 3

Do not refer to any printed material. No talking or discussion is allowed until the answer sheets are all collected.

Review:

1. The file /etc/fstab:

(A) determines which kernel to load at boot time (B) chooses filesystems the kernel will mount at boot

time (C) determines whether the kernel supports the Ext3

filesystem type (D) determines the maximum number of open files (E) lists all filesystem types supported by the kernel

2. $ ls -ld directory
drwxr-xr--   3 lee students 48 Feb 11 09:10 directory

(A) members of group staff cannot list files in

directory

(B) members of group students can change to

directory

(C) members of group students can delete files in

directory

(D) It is impossible for user lee to be a member of

group students

(E) both user lee and members of group students can delete files in directory.

3. the directory /proc contains information that is:

(A) dynamic information about the OS and hardware (B) a type of relational database; SQL is always used to

query it (C) a list of network procters

(D) is unreadable except by the OS kernel

(E) contains entirely static data that is fixed when the computer boots

Preparation:

4. A shell script:

(A) cannot support normal programming using loops and function calls

(B) is compiled into machine language before it is executed

(C) is always optimised for performance, and executes faster than Java programs

(D) is seldom used for system administration

(E) can only be written by professional programmers

5. Select one statement that is most correct:

(A) There is very little documentation for shell programming, but the O’Reilly books are good references

(B) The man page for bash does not provide enough information about the language for a reference guide

(C) There is a free online book about shell programming

(D) Shell programs are always quite long.